OpenStudy (anonymous):

Micaela and Aliyah are selling cheesecakes for a school fundraiser. Customers can buy pecan cheesecakes and chocolate marble cheesecakes. Micaela sold 8 pecan cheesecakes and 10 chocolate marble cheesecakes for a total of $240. Aliyah sold 10 pecan cheesecakes and 8 chocolate marble cheesecakes for a total of $212. Find the cost of one pecan cheesecake and one chocolate marble cheesecake.

OpenStudy (anonymous):

8x+10y=240 10x+7y=212 @phi

OpenStudy (phi):

you can simplify the first equation

OpenStudy (anonymous):

4x+5y=120

OpenStudy (phi):

ok, now you have 4x+5y=120 10x+7y=212 what's a "least common denominator" of 4 and 10 ?

OpenStudy (anonymous):

is it 2?

OpenStudy (phi):

you can practice that later. but 4*5 is 20 and 2*10 is 20

OpenStudy (phi):

the idea is to make the number in front of the x the same in both equations (I picked x because the numbers look a tiny bit simpler than for y)

OpenStudy (anonymous):

ok so does that make the number infront of x 20?

OpenStudy (phi):

if you multiply the top equation by 5 and the bottom by 2 can you do that ?

OpenStudy (anonymous):

20x+25y=600 and 20x+14y=424

OpenStudy (phi):

ok, btw, the problem says Aliyah sold 10 pecan cheesecakes and 8 chocolate marble cheesecakes for a total of $212 but you are using 10x+7y=212 ??

OpenStudy (anonymous):

oh god, I keep hitting the wrong key hahaha I meant 8!!

OpenStudy (anonymous):

I wrote it like that too, jeez I must be slow today

OpenStudy (anonymous):

let me fix this then okay. so 20x+16y=424 then

OpenStudy (phi):

ok. now you have 20x+25y=600 20x+16y=424 any ideas?

OpenStudy (anonymous):

do you subtract the bottom equation from the top?

OpenStudy (phi):

either way, but it's slightly nicer to subtract the bottom from the top

OpenStudy (anonymous):

ok so the 20x-20x cancels out and then 25y-16y=9 and 600-424=176

OpenStudy (phi):

25y-16y=9 not just 9. it is 9y

OpenStudy (anonymous):

yes 9y=176

OpenStudy (anonymous):

176 divided by 9? right? 19.555556 it says

OpenStudy (phi):

we are going to get an fraction or decimal in the answer. (It was nicer if we could use 7 instead of 8)

OpenStudy (anonymous):

haha true

OpenStudy (anonymous):

so y=19.5 and x=5.625?

OpenStudy (phi):

x and y are dollars and cents

OpenStudy (anonymous):

ok but that's right?

OpenStudy (phi):

If you don't use fractions when solving for x then you should use y with as many decimals as your calculator allows (do not round)

OpenStudy (anonymous):

well it was 19.5555556 and 5.625 I'm pretty sure

OpenStudy (phi):

you should get x= 50/9 = 5.5555...

OpenStudy (anonymous):

ahhh I must've just clicked something wrong.

OpenStudy (phi):

and at the end, round to the nearest cent: 19.56 and 5.56

OpenStudy (anonymous):

ok that makes sense. thank you so much :) I just have one question left that I've started. I'll make a new post